# A Tale of Two Imperatives: Privacy and Explainability

### v3

@@ 1 Introduction
Deep learning has achieved massive success over the past decade in several domains and high stakes Dong et al. (2021). It has been a cornerstone ever since in almost all aspects of scientific discoveries. However, it is to be noted that deep learning, although predominant in today’s scientific understandings, comes with its inherent shortcomings Talaei Khoei et al. (2023); Saeed and Omlin (2023); Boulemtafes et al. (2020). In this paper, we shall investigate the intricacies of two such pivotal shortcomings: Privacy and Explainability. We shall thereafter substantiate our findings with a famous use case from previous studies.
@@
Firstly, the deep models, although achieve superior performance across domains, are inherently prone to sensitive data leakage Boulemtafes et al. (2020). It has been extensively shown that the models can leak information about the data with which it has been trained. Even, especially in the biomedical domain, it has been shown that simple ‘linkage’ attacks can exploit anonymized electronic health records Sweeney (2015). Due to the inherently vulnerable nature of deep models, it is not only prone to severe privacy attacks including membership inference (MIA) Hu et al. (2022), model stealing Oliynyk et al. (2023), model inversion Fredrikson et al. (2015); Wang et al. (2021); Veale et al. (2018) etc but also impose a threat on the applicability of deep learning in real-world applications.
@@
Secondly, the deep models are hard to explain Saeed and Omlin (2023). Neural nets are nonlinear systems, often try to learn the (approximate) distribution of the training data. However, once a model is deployed in an open-world setting Zhu et al. (2024), there is no gold label to cross-check the prediction obtained as a result, the subtle notion of trust on the prediction is pivotal to move forward with the same. But as these models are excessively critical, how did the model come to the prediction is a non-trivial phenomenon Jentzen et al. (2023).
@@
To tackle the first problem, researchers have developed several strategies for privacy-preserving machine learning Boulemtafes et al. (2020). This includes homomorphic encryption Pulido-Gaytan et al. (2020), PATE Papernot et al. (2018), differential privacy (DP) Abadi et al. (2016) etc; depending on the setting and motivation we incorporate these methods to make the model robust against diverse privacy attacks. However, differential privacy (DP), over the years, has been established as a gold-standard for privacy-preserving machine learning Blanco-Justicia et al. (2022); Suriyakumar et al. (2021) mainly due to its worst case guarantee against any inference attacks Abadi et al. (2016); Dwork et al. (2014). Furthermore, other types of privacy attacks are not as mature and/or successful as MIA is Rigaki and Garcia (2023). In this paper, we have specifically worked on DP models [^1] Throughout the paper we use DP model(s)/explanation(s) and private model(s)/explanation(s) interchangeably. with diverse privacy guarantees aka $\epsilon$ Dwork et al. (2014).
@@
On the other hand, to tackle the explainability problem, researchers have developed an array of methods to audit the model decisions Rudin et al. (2021); Saeed and Omlin (2023). This includes (but not limited to) local Ribeiro et al. (2016); Lundberg and Lee (2017); Sundararajan et al. (2017); Smilkov et al. (2017) and global explainability methods Ibrahim et al. (2019); Lundberg et al. (2020), inherently interpretable machines Molnar et al. (2020); Ditz (2024) etc. Among these, local post-hoc explainers have emerged as a popular and widely adopted tools for model auditing in recent times Saeed and Omlin (2023). Furthermore, post-hoc explainers’ ‘plug-and-play’ nature is one of the crucial reasons for their massive success especially, in the industrial settings Bhatt et al. (2020b). In this work, we employ five popular post-hoc explainers relevant to our use case to obtain explanations.
@@
Although privacy and explainability aspects have been vastly explored independently, there is little to no work incorporating them together. Despite the rapid advancements in AI, integrating privacy and explainability in high-stakes domains remains an unsolved and pressing challenge. Existing research has treated privacy and explainability as separate challenges, leaving a significant gap in understanding their interplay. This gap becomes critical in high-stakes applications where both are non-negotiable. Researchers in privacy-preserving machine learning have investigated several aspects of differential privacy such as the scalability of DP models Beltran et al. (2024), their fairness Fioretto et al. (2022), robustness Tursynbek et al. (2020), privacy-utility trade-off Zheng et al. (2024) etc. On the other side, researchers have worked on different aspects of explainable AI (XAI) such as faithfulness Lyu et al. (2024), robustness Sinha et al. (2021); Ivankay et al. (2022), and the quality of explanations Zhou et al. (2021) to name a few. In this paper, we take the first comprehensive step toward bridging the divide between privacy-preserving models and explainable AI in high-stakes domains. We explore the unique challenges that emerge when attempting to integrate privacy and explainability aspects in high-stakes applications. We investigate the underlying causes of these challenges, examine the trade-offs involved, and discuss key considerations necessary for developing frameworks to successfully incorporate these two critical aspects effectively.
@@
Integrating privacy and explainability is non-trivial, but addressing these issues is crucial for ensuring trustworthy and effective AI systems in high-stakes. In this context, we argue that it is equally important to identify use cases where achieving ‘the best of both worlds’ is not just desirable but a requirement. For example, if the training data is open source or an interpretable model (e.g. decision trees) is considered for the study, our research may not be well-motivated or the findings are not ‘worthy’. This is why, for this paper, we have considered the well-established use-case of disease detection from chest x-ray Al-qaness et al. (2024). The reason to choose the same is that medical records are (almost) always subject to privacy preservation and it is of paramount importance that, with explanations, the stakeholders including both physician and patients can trust the model predictions. Post-hoc explainers, which we use in this paper, have been extensively utilized for our chosen and similar use-cases in past Gaudio (2023); E. Ihongbe et al. (2024); Gwinner et al. (2024); Saxena et al. (2022); Ifty et al. (2024); prasad Koyyada and Singh (2023); Sarp et al. (2023).
@@
The dual mandates of the Right-to-Privacy (RTP) Thomson (1975) and the Right-to-Explanation (RTE) Vredenburgh (2022), increasingly imposed by governments for high-stakes AI applications, underscore the urgency of addressing these intertwined imperatives. Despite their critical importance, prior research lacks a comprehensive and systematic feasibility study that explores the practical integration of these rights within AI systems. Fundamental questions remain unanswered: How should we evaluate the quality of explanations for DP models? Can existing popular local post-hoc explainers reliably function in this setting, or do they falter under the constraints of differential privacy? If they fail, what alternative methods can generate comparable private explanations? Our work makes a fundamental contribution by systematically addressing these pressing questions, providing not only foundational clarity but also actionable insights grounded in rigorous empirical analysis. Formally, our central research question in this paper is:
@@
  • We investigate the interplay between the DP models and post-hoc explainers subject to the desiderata proposed;
@@
The rest of the paper is organized as follows. We brief the preliminary concepts of DP and post-hoc explainers. We point out the ways an adversary can attack, followed by an extensive discussion on several quantifiable notions for explanations’ quality, and their inherent shortcomings in Section 3. We formally introduce our proposed desiderata and measures to quantify explanation quality in Section 4. We detail our experimental setup in Section 5. We discuss and our findings and mechanistically interpret the models to explain our findings in Sections 6 and 7. Next, we explore Local Differential Privacy and its applicability in our context, in Section 8. We outline our privacy-preserving pipeline in Section 9. We present the related works in Section 10. Finally, we point out the limitations of our work in Section 11, and conclude in Section 12.
@@ 2 Preliminaries
Privacy-preserving machine learning is the building block for RTP in modern day ML systems. Researchers have developed numerous strategies for privacy-preserving machine learning, including homomorphic encryption, PATE, and differential privacy (DP) Boulemtafes et al. (2020); Papernot et al. (2018); Abadi et al. (2016). Among these, DP has emerged as the gold standard due to its strong worst-case guarantee against inference attacks Blanco-Justicia et al. (2022); Dwork et al. (2014). In this work, we focus on DP models with varying privacy guarantees.
@@
DP comes with a few intersting properties such as sequential composition, parallel composition and, post-processing. In our study, post-processing is most relevant which let the user perform arbitrary operations on the output of a DP mechanism Dwork et al. (2014).
To make a model differentially private (DP), it undergoes privacy-preserving training, with DP-SGD being the most prominent method Ponomareva et al. (2023). It introduces Gradient clipping ensuring bounded contributions of individual data points and noise to the gradients during each update to ensure differential privacy Abadi et al. (2016). Gaussian noise is commonly used due to its well-analyzed privacy guarantees. Alternatively, laplacian noise is employed in some contexts where bounded sensitivity is easier to calculate Dwork et al. (2014). For DP-SGD, gaussian noise is predominantly used whereas, Laplacian noise is popular in Local Differential Privacy (LDP) Dwork et al. (2014); Ponomareva et al. (2023).
@@
Explainable machine learning is crucial for Right-to-Explainability (RTE). As mentioned earlier, this paper focuses on local, post-hoc explainers (denoted as $\mathcal{I}$) due to their “plug-and-play” nature, which makes them highly practical, especially for industrial applications Bhatt et al. (2020b). Broadly, these explainers can be classified into two categories: Perturbation-based methods and Gradient-based methods. Both types of methods take a model ($g$) and a datapoint ($x$) as input and output a feature attribution score(s) against $g(x)$. A a feature attribution score (FAS) (or a feature attribution vector) consists of scores assigned to individual features of the given input representing their importance towards the classification the given model comes up with. A positve attribution score implies a feature has positively contributed towards the classification and negative attribution score shows that a feature does not positively contribute towards the classification. Given an explainer and a datapoint ($x$), we denote the FAS which explains the prediction of $\mathcal{M}$ and $\mathcal{M^{\prime}}$ as s and $s^{\prime}$ respectively; given $\mathcal{M}(x)=$ $\mathcal{M^{\prime}}(x)$.
@@ 3.1 The Privacy Aspect
From the point of $\mathbf{A}$, we identify two distinct ways of attacking and leaking the information from the trained model [^2] As in this paper we are working with DP models, we have considered MIA as the only type of potential privacy attack for simplicity. However, it has been shown that other classes of privacy attacks (model stealing, model inversion etc) are also very much possible with the two ways of attacking mentioned and DP potentially can safeguard against a few other attacks besides MIA Rigaki and Garcia (2023).:
@@
    For a given data-point $x$, assuming $\mathbf{A}$ has access to the output vector V from the trained model $\mathcal{M}$, $\mathbf{A}$ can perform an MIA on the model Hu et al. (2022).
@@
    Assuming $\mathbf{A}$ has access to the feature attribution score s of $x$ generated by $\mathcal{I}$ subject to $\mathcal{M}(x)$, $\mathbf{A}$ can leverage s to execute an MIA on $\mathcal{M}$ Shokri et al. (2021), especially when $\mathcal{I}$ is faithful.
However, as $\mathcal{M^{\prime}}$, is the DP counterpart of $\mathcal{M}$, it inherits the post-processing property Dwork et al. (2014) which makes any mechanism including obtaining V, s, applied over $\mathcal{M^{\prime}}$ to produce results which are also DP, so we can mitigate both of the breaches.
@@ 3.2 The Explainability Aspect
As discussed earlier, in this study we are exclusively considering local post-hoc explainers Lundberg et al. (2020); Huber et al. (2021); Lundberg and Lee (2017). The quality these local explanations are judged across several parameters Hedström et al. (2023): Faithfulness Miró-Nicolau et al. (2024); Lyu et al. (2024), Robustness Alvarez-Melis and Jaakkola (2018); Mishra et al. (2021), Localization Zhang et al. (2018); Theiner et al. (2022); Kohlbrenner et al. (2020); Arias-Duart et al. (2022), Complexity Chalasani et al. (2020); Bhatt et al. (2020a); Nguyen and Martínez (2020), Randomization Adebayo et al. (2018); Sixt et al. (2020), are mostly accentuated in previous studies.
Unequivocally, faithfulness is the most crucial among all Miró-Nicolau et al. (2024); Lyu et al. (2024). Faithfulness is loosely defined as how well the explainer reflects the underlying reasoning of the model and has been extensively quantified in diverse ways Li et al. (2023); Hedström et al. (2023): Sufficiency (SF) Dasgupta et al. (2022), Infidelity (IF) Yeh et al. (2019), Insertion/Detection AUC (I/D-AUC) Petsiuk et al. (2018), Pixel Flipping (PF) Bach et al. (2015), Iterative Removal of Features (IROF) Rieger and Hansen (2020), Ordered Perturbation based metrics (OPs) Samek et al. (2016) are to mention a few. However, the evaluation metric(s) are not necessarily flawless; which we’re going to discuss next.
@@ 3.3 Pitfalls of the Evaluation Metrics
Firstly, in the perturbation-based metrics such as PF, I/D-AUC, IROF, IF etc while applying operations such as flipping the pixels, inserting and/or deleting features, performing meaningful perturbation on the input space to generate synthetic inputs as a part of their evaluation process, do not crosscheck whether the generated input is out-of-distribution with respect to the trained model Hase et al. (2021); Chang et al. (2018). Secondly, metrics such as OPs are often excessively similar to the mechanics of the explainer itself. Instead of evaluating faithfulness, these methods primarily compute the similarity between the evaluation metric and explanation techniques, assuming the evaluation metric itself to be the ground truth Ju et al. (2021). Li et al. Li et al. (2023) has acknowledged the same in their benchmark $\mathcal{M}^{4}$ that LIME Ribeiro et al. (2016) and OPs are methodologically similar thus, the evaluation maybe skewed. Thirdly, previous studies have extensively shown that test-time input ablation is often prone not only to generating out-of-distribution (OOD) synthetic input Hase et al. (2021); Haug et al. (2021); Chang et al. (2018); Janzing et al. (2020) but also are socially misaligned Jacovi and Goldberg (2021). Thus, metrics for example I/D-AUC, IROF etc are suspected to be severely misleading. Quantitative metrics such as SF are also substantially constrained and do not provide a universal overview of faithfulness. Lastly, all these evaluations are based on naive assumptions (e.g.: erasure Jacovi and Goldberg (2020)), derived from a set of seemingly valid observations. Therefore, these quantitative metrics are not necessarily axiomatically valid. Even axiomatic necessary tests such as Model Parameter Randomization test Adebayo et al. (2018) happen to have a set of empirical confounders Yona and Greenfeld (2021); Kokhlikyan et al. (2021); Bora et al. (2024). To the best of our knowledge, there has not been any universally valid necessary and sufficient approach for faithfulness evaluation Lyu et al. (2024).
Furthermore, the lack of a unified faithfulness evaluation method and the absence of ground truth for explainers lead to the well-known disagreement problem Evans et al. (2022); Neely et al. (2021); Roy et al. (2022); Krishna et al. (2022). Even if explainers are faithful, their inherent mechanisms for calculating feature importance can lead to different explanations. For instance, Han et al. Han et al. (2022) demonstrated that while a subset of commonly used XAI methods are all local function approximators, they employ different kernel and loss functions. Consequently, like faithfulness, other parameters for evaluating explainers are also contentious. For example, adversarial robustness (of the explainer) assumes that similar inputs with similar outputs should yield similar explanations Sinha et al. (2021); Ivankay et al. (2022). However, Ju et al. Ju et al. (2021) has empirically shown that the change in attribution scores may be because the model’s reasoning process has genuinely changed, rather than because the attribution method is unreliable. Moreover, this assumption is mainly valid when the model is astute Bhattacharjee and Chaudhuri (2020) and doesn’t necessarily apply to explainers that don’t perform local function approximation for feature importance estimation Han et al. (2022). In the subsequent section(s) we’ll try to figure out whether we can bypass these limitations for our study.

### v4

# Reconciling Privacy and Explainability in High-Stakes: A Systematic Inquiry

@@ 1 Introduction
Deep learning has achieved massive success over the past decade in several domains and high stakes (Dong et al., 2021). It has been a cornerstone ever since in almost all aspects of scientific discoveries. However, it is to be noted that deep learning, although predominant in today’s scientific understandings, comes with its inherent shortcomings (Talaei Khoei et al., 2023; Saeed and Omlin, 2023; Boulemtafes et al., 2020). In this paper, we shall investigate the intricacies of two such pivotal shortcomings: Privacy and Explainability. We shall thereafter substantiate our findings with a famous use case from previous studies.
@@
Firstly, the deep models, although achieve superior performance across domains, are inherently prone to sensitive data leakage (Boulemtafes et al., 2020). It has been extensively shown that the models can leak information about the data with which it has been trained. Even, especially in the biomedical domain, it has been shown that simple ‘linkage’ attacks can exploit anonymized electronic health records (Sweeney, 2015). Due to the inherently vulnerable nature of deep models, it is not only prone to severe privacy attacks including membership inference (MIA) (Shokri et al., 2017; Hu et al., 2022), model stealing (Oliynyk et al., 2023), model inversion (Fredrikson et al., 2015; Wang et al., 2021; Veale et al., 2018) etc but also impose a threat on the applicability of deep learning in real-world applications.
@@
Secondly, the deep models are hard to explain (Saeed and Omlin, 2023). Neural nets are nonlinear systems, often try to learn the (approximate) distribution of the training data. However, once a model is deployed in an open-world setting (Zhu et al., 2024), there is no gold label to cross-check the prediction obtained as a result, the subtle notion of trust on the prediction is pivotal to move forward with the same. But as these models are excessively critical, how did the model come to the prediction is a non-trivial phenomenon (Jentzen et al., 2023).
@@
To tackle the first problem, researchers have developed several strategies for privacy-preserving machine learning (Boulemtafes et al., 2020). This includes homomorphic encryption (Pulido-Gaytan et al., 2020), PATE (Papernot et al., 2018), differential privacy (DP) (Abadi et al., 2016) etc; depending on the setting and motivation we incorporate these methods to make the model robust against diverse privacy attacks. However, differential privacy (DP), over the years, has been established as a gold-standard for privacy-preserving machine learning (Blanco-Justicia et al., 2022; Suriyakumar et al., 2021) mainly due to its worst case guarantee against any inference attacks (Abadi et al., 2016; Dwork et al., 2014). Furthermore, other types of privacy attacks are not as mature and/or successful as MIA is (Rigaki and Garcia, 2023). In this paper, we have specifically worked on DP models [^1] Throughout the paper we use DP model(s)/explanation(s) and private model(s)/explanation(s) interchangeably. with diverse privacy guarantees aka $\epsilon$ (Dwork et al., 2014).
@@
On the other hand, to tackle the explainability problem, researchers have developed an array of methods to audit the model decisions. This includes (but not limited to) local and global explainability methods, inherently interpretable machines, etc (Saeed and Omlin, 2023). Among these, local post-hoc explainers have emerged as popular and widely adopted tools for model auditing in recent times due to their ‘plug-and-play’ nature (Bhatt et al., 2020). In this work, we employ five popular post-hoc explainers relevant to our use case to obtain explanations.
@@
Although privacy and explainability aspects have been vastly explored independently, there is little to no work incorporating them together. Despite the rapid advancements in AI, integrating privacy and explainability in high-stakes domains remains an unsolved and pressing challenge. Existing research has treated privacy and explainability as separate challenges, leaving a significant gap in understanding their interplay. This gap becomes critical in high-stakes applications where both are non-negotiable. Researchers in privacy-preserving machine learning have investigated several aspects of differential privacy such as the scalability of DP models (Beltran et al., 2024), their fairness (Fioretto et al., 2022), robustness (Tursynbek et al., 2020), privacy-utility trade-off (Zheng et al., 2024) etc. On the other side, researchers have worked on different aspects of explainable AI (XAI) such as faithfulness (Lyu et al., 2024), robustness (Alvarez-Melis and Jaakkola, 2018), and the quality of explanations (Zhou et al., 2021) to name a few. In this paper, we take the first comprehensive step toward bridging the divide between privacy-preserving models and explainable AI in high-stakes domains. We explore the unique challenges that emerge when attempting to integrate privacy and explainability aspects in high-stakes applications. We investigate the underlying causes of these challenges, examine the trade-offs involved, and discuss key considerations necessary for developing frameworks to successfully incorporate these two critical aspects effectively.
@@
Integrating privacy and explainability is non-trivial, but addressing these issues is crucial for ensuring trustworthy and effective AI systems in high-stakes. In this context, we argue that it is equally important to identify use cases where achieving ‘the best of both worlds’ is not just desirable but a requirement. For example, if the training data is open source or an interpretable model (e.g. decision trees) is considered for the study, our research may not be well-motivated or the findings are not ‘worthy’. This is why, for this paper, we have considered the well-established use-case of disease detection from chest x-ray (Al-qaness et al., 2024). The reason to choose the same is that medical records are (almost) always subject to privacy preservation and it is of paramount importance that, with explanations, the stakeholders including both physician and patients can trust the model predictions. Post-hoc explainers, which we use in this paper, have been extensively utilized for our chosen and similar use cases in past (E. Ihongbe et al., 2024; Saxena et al., 2022; Ifty et al., 2024).
@@
The dual mandates of the Right-to-Privacy (RTP) (Thomson, 1975) and the Right-to-Explanation (RTE) (Vredenburgh, 2022), increasingly imposed by governments for high-stakes AI applications, underscore the urgency of addressing these intertwined imperatives. Despite their critical importance, prior research lacks a comprehensive and systematic feasibility study that explores the practical integration of these rights within AI systems. Fundamental questions remain unanswered: How should we evaluate the quality of explanations for DP models? Can existing popular local post-hoc explainers reliably function in this setting, or do they falter under the constraints of differential privacy? If they fail, what alternative methods can generate comparable private explanations? Our work makes a fundamental contribution by systematically addressing these pressing questions, providing not only foundational clarity but also actionable insights grounded in rigorous empirical analysis. Formally, our central research question in this paper is:
@@
  • We investigate the interplay between the DP models and post-hoc explainers, subject to the desiderata proposed;
@@
The rest of the paper is organized as follows. We brief the preliminary concepts of DP and post-hoc explainers in Section 2. We point out the potential privacy breaches, followed by an extensive discussion on several quantifiable notions for explanations’ quality, and their inherent shortcomings in Section 3. We formally introduce our proposed postulate and quantifiable measures for the same in Section 4. We detail our experimental setup in Section 5. We discuss our findings in Section 6. We mechanistically interpret the models to explain our findings in Section 7. Next, we explore Local Differential Privacy and its applicability in our context, in Section 8. We outline our novel software pipeline to reconcile privacy and explainability in Section 9. We present the related works in Section 10. Finally, we conclude in Section 11.
@@ 2 Preliminaries
Privacy-preserving machine learning is the building block for RTP in modern day ML systems. Researchers have developed numerous strategies for privacy-preserving machine learning, including homomorphic encryption, PATE, and differential privacy (DP) (Boulemtafes et al., 2020; Papernot et al., 2018; Abadi et al., 2016). Among these, DP has emerged as the gold standard due to its strong worst-case guarantee against inference attacks (Blanco-Justicia et al., 2022; Dwork et al., 2014). In this work, we focus on DP models with varying privacy guarantees.
@@
DP comes with a few interesting properties such as sequential composition, parallel composition and, post-processing. In our study, post-processing is most relevant which lets the user perform arbitrary operations on the output of a DP mechanism without hampering its privacy (Dwork et al., 2014).
To make a model differentially private (DP), it undergoes privacy-preserving training, with DP-SGD being the most prominent method (Ponomareva et al., 2023). It introduces Gradient clipping ensuring bounded contributions of individual data points and noise to the gradients during each update to ensure differential privacy (Abadi et al., 2016). Gaussian noise is commonly used due to its well-analyzed privacy guarantees. Alternatively, laplacian noise is employed in some contexts where bounded sensitivity is easier to calculate (Dwork et al., 2014). For DP-SGD, gaussian noise is predominantly used whereas, Laplacian noise is popular in Local Differential Privacy (LDP) (Dwork et al., 2014; Ponomareva et al., 2023).
@@
Explainable machine learning is crucial for Right-to-Explainability (RTE). As mentioned earlier, this paper focuses on local, post-hoc explainers (denoted as $\mathcal{I}$). Broadly, these explainers can be classified into two categories: Perturbation-based methods and Gradient-based methods. Both types of methods take a model ($g$) and a datapoint ($x$) as input and output a feature attribution score(s) against $g(x)$. A feature attribution score (FAS) (or a feature attribution vector) consists of scores assigned to individual features of the given input representing their importance towards the classification the given model comes up with. A positive attribution score implies a feature has positively contributed towards the classification and a negative attribution score shows that a feature does not positively contribute towards the classification. Given an explainer and a datapoint ($x$), we denote the FAS which explains the prediction of $\mathcal{M}$ and $\mathcal{M^{\prime}}$ as s and $s^{\prime}$ respectively; given $\mathcal{M}(x)=$ $\mathcal{M^{\prime}}(x)$.
@@ 3.1 The Privacy Aspect
From the point of $\mathbf{A}$, we identify two distinct ways of attacking and leaking the information from the trained model [^2] As in this paper we are working with DP models, we have considered MIA as the only type of potential privacy attack for simplicity. However, it has been shown that other classes of privacy attacks (model stealing, model inversion, etc) are also very much possible with the two ways of attacking mentioned, and DP potentially can safeguard against a few other attacks besides MIA (Rigaki and Garcia, 2023). However, these fall beyond the scope of the present study.:
@@
    For a given data-point $x$, assuming $\mathbf{A}$ has access to the output vector V from the trained model $\mathcal{M}$, $\mathbf{A}$ can perform an MIA on the model (Hu et al., 2022).
@@
    Assuming $\mathbf{A}$ has access to the feature attribution score s of $x$ generated by $\mathcal{I}$ subject to $\mathcal{M}(x)$, $\mathbf{A}$ can leverage s to execute an MIA on $\mathcal{M}$ (Shokri et al., 2021), especially when $\mathcal{I}$ is faithful.
However, as $\mathcal{M^{\prime}}$ is the DP counterpart of $\mathcal{M}$, it inherits the post-processing property (Dwork et al., 2014) which makes any mechanism including obtaining V, $s^{\prime}$, from $\mathcal{M^{\prime}}$ to produce results which are also DP, i.e., any post-hoc explanations for an ($\epsilon,\delta$)-DP model also satisfy ($\epsilon,\delta$)-DP for the training dataset (Patel et al., 2022). Thus, using $\mathcal{M^{\prime}}$ we can mitigate both of the breaches.
@@ 3.2 The Explainability Aspect
As discussed earlier, in this study, we are exclusively considering local post-hoc explainers (Lundberg et al., 2020; Huber et al., 2021; Lundberg and Lee, 2017). The quality these local explanations are judged across several parameters (Hedström et al., 2023): Faithfulness (Lyu et al., 2024), Robustness (Alvarez-Melis and Jaakkola, 2018), Localization (Zhang et al., 2018; Theiner et al., 2022), Complexity (Chalasani et al., 2020; Nguyen and Martínez, 2020), Randomization (Adebayo et al., 2018), are mostly accentuated in previous studies.
Unequivocally, faithfulness is the most crucial among all (Lyu et al., 2024). Faithfulness is loosely defined as how well the explainer reflects the underlying reasoning of the model and has been extensively quantified in diverse ways (Li et al., 2023; Hedström et al., 2023): Sufficiency (SF) (Dasgupta et al., 2022), Infidelity (IF) (Yeh et al., 2019), Insertion/Detection AUC (I/D-AUC) (Petsiuk et al., 2018), Pixel Flipping (PF) (Bach et al., 2015), Iterative Removal of Features (IROF) (Rieger and Hansen, 2020), Ordered Perturbation based metrics (OPs) (Samek et al., 2016) are to mention a few. However, the evaluation metric(s) are not necessarily flawless, which we’re going to discuss next.
@@ 3.3 Pitfalls of the Evaluation Metrics
Firstly, in the perturbation-based metrics such as PF, I/D-AUC, IROF, IF etc while applying operations such as flipping the pixels, inserting and/or deleting features, performing meaningful perturbation on the input space to generate synthetic inputs as a part of their evaluation process, do not crosscheck whether the generated input is out-of-distribution with respect to the trained model (Hase et al., 2021; Chang et al., 2018). Secondly, previous studies have extensively shown that test-time input ablation is often prone not only to generating out-of-distribution (OOD) synthetic input (Hase et al., 2021; Haug et al., 2021) but also are socially misaligned (Jacovi and Goldberg, 2021). Thus, metrics for example I/D-AUC, IROF etc are suspected to be severely misleading; even metrics such as SF are also substantially constrained and do not provide a universal overview of faithfulness. Thirdly, metrics such as OPs are often excessively similar to the mechanics of the explainer itself. Instead of evaluating faithfulness, these methods primarily compute the similarity between the evaluation metric and explanation techniques, assuming the evaluation metric itself to be the ground truth (Ju et al., 2021). Li et al. (Li et al., 2023) has acknowledged the same in their benchmark $\mathcal{M}^{4}$ that LIME (Ribeiro et al., 2016) and OPs are methodologically similar thus, the evaluation maybe skewed. Lastly, all these evaluations are based on naive assumptions (e.g.: erasure (Jacovi and Goldberg, 2020)), derived from a set of seemingly valid observations. Therefore, these quantitative metrics are not necessarily axiomatically valid. Even axiomatic necessary tests such as Model Parameter Randomization test (Adebayo et al., 2018) happen to have a set of empirical confounders (Yona and Greenfeld, 2021; Kokhlikyan et al., 2021; Bora et al., 2024). To the best of our knowledge, there has not been any universally valid necessary and sufficient approach for faithfulness evaluation (Lyu et al., 2024).
Furthermore, the absence of ground truth for models’ reasoning makes it an open challenge to universally quantify faithfulness and evaluate a faithfulness measure precisely, i.e. when a universal ground truth for reasoning is not available, axiomatically quantifying how much a faithfulness measure is more reliable than others is challenging; however, in the existing literature, faithfulness measures evolve by addressing the discovered shortcomings of their predecessors (Lyu et al., 2024). This also leads to the well-known disagreement problem (Krishna et al., 2022): even if explainers are faithful, their inherent mechanisms for calculating feature importance can lead to different explanations. In the subsequent section(s), we’ll try to figure out whether we can bypass these limitations for our study.